# Integrated modelling of equilibrium and transport in axisymmetric magnetic mirror fusion devices

### v1

@@ 4.1 Assumptions and constraints
Beginning with the modelling assumptions and fixed parameters; two large physics assumptions were made in this study to overcome the limitations of present modelling tools. First, it was assumed that the end plug, for the purposes of equilibrium and transport modelling, could be treated as a standalone simple mirror. Second, the question of MHD stabilization of the end plug was ignored. The justification of the first assumption lies in the fact that the ion populations in the plug and the central cell are to lowest order "detached". Ions confined in the central cell do not enter the plugs and ions confined in the plugs do not enter the central cell. Similarly ions that are not confined in the plug are also not confined in the central cell and vice-versa. This property of tandem mirrors ends up being useful for reasons that will be discussed momentarily. In fusion-relevant conditions, confinement is good in both the plugs and the central cell and the unconfined population of ions flowing through either region is small, acting as a (relatively) low density stream which passes outwards until it reaches the expander region (Cohen et al., 1980). However, in a classical tandem mirror the electron population is able to transit between the plug and central cell, and alpha particle heating from the central cell is sufficient to enable us to reach high electron temperatures. In fact, the predicted electron temperatures in a tandem are higher than they otherwise would be in a standalone end plug. This makes the estimates of plug performance in this model conservative with respect to electron drag, as the increasing electron temperature will decrease electron drag driven ion losses and improve ion confinement (however there can be increased ion losses due to larger ambipolar potentials which may incentivize using higher neutral beam energies). A more detailed analysis of how varying $T_{e}$ affects ion confinement can be found in Appendix A. For the purposes of the simulations here, it was assumed that the plasma is MHD stable and whichever actuator used to ensure MHD stability was assumed to not substantially perturb equilibrium transport. This topic will be tackled in more detail in future work.
@@
As mentioned earlier, the detached nature of the end plug and central cell ion distributions can be used to improve performance. In a tandem, the end plug ion composition need not be the same as the central cell as the end plug and central cell ion populations are not exchanged (Fowler & Logan, 1977; Cohen et al., 1980). This presents a unique opportunity to enhance end plug physics performance and simplify end plug engineering. To this end, end plugs fueled with only tritium are used in conjunction with a deuterium and tritium fueled central cell. This combination maximizes end plug performance and reduces the nuclear engineering challenges presented by the end plug. The physical argument for this operational mode is as follows: ion confinement in the end plug is dominated by two processes: ion-ion scattering and electron drag. The particle confinement times associated with these two processes are (Logan et al., 1980):
@@
At typical tandem operational parameters, confinement is assumed to be electron drag dominated, $\tau_{d}\ll\tau_{s}$ (Logan et al., 1980; Mirin et al., 1983; Hua & Fowler, 2004; Fowler et al., 2017). As the end plug ion population composition need not be the same as the central cell, performance can be enhanced by choosing ions with large $\mu/Z^{2}$. In addition, choosing an ion mix to suppress fusion neutron rates in the end plug can substantially simplify engineering by reducing radiation shielding requirements for components such as magnets and NBI. Suppressing fusion rates in the end plugs is acceptable as the total fusion power produced by the end plug is a small component of the overall fusion power in a tandem, since the end plug volume is much less than central cell volume, $V_{p}\ll V_{c}$. If taken to its logical conclusion, this argument suggests that a tritium (T) only end plug is optimal with respect to classical transport and neutron damage. Tritium has the highest $\mu/Z^{2}$ ratio of any standard fusion fuel ion, and the T-T fusion cross-section is nearly an order of magnitude lower than the D-D fusion cross section, and T-T fusion produces neutrons with energies of $\sim 3~{}\mathrm{MeV}$ (Brown et al., 2018). Thus, a T-only end plug provides a factor of $\sim 1.2$ increase in confinement performance and a many-order of magnitude reduction in the neutron rate as well as a significant reduction in neutron energy relative to a D-T end plug (T fueled plugs will have inferior microstability to D or H plugs as the T Larmor radius is larger, making it easier to excite diamagnetic drift waves. However, it will be shown that it should be possible to avoid diamagnetic drift waves with large $a/\rho_{i}$ even with T plugs).
@@
Another assumption is the use of only negative-ion neutral beams for end plug heating in steady-state (a discussion about the limitations associated with high-harmonic RF heating coupled with positive-ion NBI, similar to what is planned in WHAM, and the reason it is not used in this work is discussed in Appendix B). The use of ECH has also been limited to achieving breakdown (assuming a scheme like that found in Yakovlev et al. (2016)). Presently, the use of any significant quantity of ECH in CQL3D-m causes the potential profile to invert. This behavior is actually promising as it indicates that it may be possible to generate thermal barriers (Baldwin & Logan, 1979), but at the moment it causes significant problems with the bounce-averaging and ambipolar potential calculations in CQL3D-m. Thus, further investigation of ECH is left for future work. Neutral transport into the tandem mirror plasma has also been neglected. This effect is anticipated to be of some importance in small relatively low density mirrors such as WHAM, where neutrals recycled from the edge can enter the plasma and charge exchange with well-confined beam ions causing them to be lost and replacing them with a promptly lost cold ion. However, fusion-grade plasmas like those under consideration here are opaque to edge neutrals, and the effect of these neutrals on transport can be ignored. The final assumption made here is that transport in the end plug will remain classical during operation. To this end, it has been insured that the DCLC stability condition is well satisfied in the plasmas optimized here.
@@
In addition to physics assumptions, some engineering constraints were imposed on the mirror design here. First, the field of the two large bore high-field mirror coils placed at either end of the end plug (denoted HF) was fixed at $25~{}\mathrm{T}$. Feasibility studies of HTS coils at these fields have found that such designs should be viable at bores sizes appropriate for fusion applications, $\sim 1$ m warm bore diameter. A single low-field coil (denoted LF) was placed at the center of the end plug to give control of the central field. The magnetic configuration for an example case is shown in Figure 6. The minimum length $\ell$ (where subscript $p$ on end plug quantities in this section has been dropped for simplicity), defined as the mirror throat to mirror throat distance. This limit was chosen not by the finite Larmor radius (FLR) stabilization condition for $m\geq 2$ modes, but rather by engineering constraints. The substantial size of the neutral curvature central cell in addition to the good curvature expanders is expected to provide significant stabilization against high $m$ modes (Roberts & Taylor, 1962; Rosenbluth et al., 1962; Rosenbluth, 1965; Pearlstein, 1966; Anikeev et al., 1992; Ryutov et al., 2011; Fowler et al., 2017). Active stabilization schemes are anticipated to provide for remaining MHD stabilization needs. Here, $\ell\geq 4.5~{}\mathrm{m}$ has been chosen as with values of $\ell<4.5~{}\mathrm{m}$ require large reverse polarity currents in the LF coil to cancel out the HF field. Additionally, very short end plugs will almost certainly impede NBI port access and require the use of inefficient NBI launch angles that limit beam path length, causing reduced NBI absorption. The optimal NBI configuration for obtaining sloshing ion distributions, maximizing beam path length, and maintaining good confinement (noting there is an effective mirror ratio reduction associated with shallower NBI launch angles) occurs at, or very close to, a neutral beam with $\theta_{\mathrm{NBI}}=45^{\circ}$ targeted at the midplane. To maintain an effective $45^{\circ}$ ion velocity-space angle and obtain a sloshing ions distribution for different beam launch angles, the beam’s target location must be moved to different values of $B/B_{0}$. The reduction factor in beam path length incurred by this move is in a simple mirror is $\sqrt{(1+\cot^{2}\theta_{\mathrm{NBI}})/4}/\sin\theta_{\mathrm{NBI}}$. To simplify this study, only the near optimal case of $\theta_{\mathrm{NBI}}=45^{\circ}$ with the beam directed at the midplane was examined in detail (the NBI here used a $0.4~{}\mathrm{m}$ by $0.8~{}\mathrm{m}$ source with no beam divergence as a placeholder until more realistic NBI specifications are known). In later work, with better defined, beam, magnet, and radiation shielding parameters, this angle may be slightly adjusted to prevent beam-line interference with components.
@@
DCLC $\underline{\text{DCLC Stability}}$
$\frac{a_{0}}{\rho_{i}}\geq 25$ (33a)
@@
for $a_{m}$ and $B_{0}$, where (33a) is a slightly modified version of the the drift-cyclotron wave stability criterion which appears in Baldwin (1977) (if the drift-cyclotron instability is stable the DCLC will not grow), and (33b) uses the analytic NBI stopping cross sections $\sigma_{s}$ from Janev et al. (1989). Using the assumptions outlined above, (33) may be rewritten (taking (33a) to be equal to 25):
@@
From the POPCON analysis performed in Section 3, it was identified that plasma radius $a_{m}\gtrsim 0.15~{}\mathrm{m}$ and density $\langle n\rangle\sim$1.0-1.5×$\times 10^{20}~{}\mathrm{m}^{-3}$. Plotting contours of $P_{\mathrm{NBI}}$ and $a_{m}$ versus $E_{\mathrm{NBI}}$ and $\langle n\rangle$ produces a plot like that shown in Figure 7. This plot indicates an operating region accessible with reasonable $P_{\mathrm{NBI}}$ values $<25~{}\mathrm{MW}$ exists between $E_{\mathrm{NBI}}$ of $200~{}\mathrm{keV}$ and $300~{}\mathrm{keV}$. However, the method used to constrain NBI absorption with (33b) is perhaps suboptimal. It has not accounted for that, despite not being absorbed as readily at high energies, the beam may still be more efficient at fueling, and simplified scaling laws were used here rather than an actual Fokker-Planck calculation. This will be corrected in the next part of this section with more detailed simulations using the RealTwin.
4.2 Parametric simulations of end plug performance
Parameter Range
$a_{m}$ 0.15 - 0.25 m
$\ell_{p}$ 4.5 m
$B_{m}$ 25 T
$B_{0}$ 4.0 - 6.0 T
$E_{\mathrm{NBI}}$ 240 - 480 keV
$P_{\mathrm{NBI}}$ 10 - 20 MW
Table 2: End plug parameters used in the parametric scan
With the constraints in the previous section established, a parametric scan in $B_{0}$, $P_{\mathrm{NBI}}$, $E_{\mathrm{NBI}}$ and $a_{m}$ was carried out with the RealTwin model described in Section 2 was specified. This scan used the end plug parameters shown in Table 2. For this work, points in the scan were selected over a linear range for each of the parameters. This scoping method is coarse, but substantial constraints were already placed on the problem through the use of the POPCON analysis in Section 3 and the analytic arguments used to constrain the physics and engineering earlier in this section. This made the coarse high fidelity scan sufficient to deliver an optimized plug that exceeded the requirements identified by the POPCON analysis. A scan in $\ell$ was also performed to determine if scaling power density with $\ell$ versus $P_{\mathrm{NBI}}$ was substantially different (changing the aspect ratio of the plug can slightly change NBI path length and modify NBI absorption). Varying $\ell$, however, had effectively the same impact on power density and plug performance as varying $P_{\mathrm{NBI}}$. In the rest of the discussion here will focus on the highest performing $\ell=4.5~{}\mathrm{m}$ case.
The CQL3D-m simulations performed in this section used 24 flux surfaces evenly spaced from 0.01 to 0.9 in normalized square-root poloidal flux $\sqrt{\psi_{n}}$, and the axial $z$ grid used to formulate the Pastukhov response along the field line had 64 evenly spaced points extending from the midplane to the mirror throat. The distribution functions were discretized over 500 points in momentum-per-rest-mass $u_{0}$ with a grid which used increased resolution at low energies as to accurate capture the ion distributions and 400 evenly spaced points in the pitch angle $\vartheta$. The maximum energy in the momentum-per-rest-mass grid was set to correspond to a $250~{}\mathrm{keV}$ electron energy. The CQL3D-m simulations used 800 timesteps. The time advance was initially set to $1~{}\mathrm{ms}$ for the first 200 time steps, then increased to $2~{}\mathrm{ms}$ for the next 200 time steps, and then further increased to $4~{}\mathrm{ms}$ for the remaining 400 timesteps providing an overall simulation runtime of $2.2~{}\mathrm{s}$, several $\tau_{p}$. Ramped time-stepping was used to maintain numerical stability, as despite being an implicit time-advance code, the iterations of CQL3D-m with Freya and Pleiades and the internal Pastukhov potential solve are explicit. The Freya NBI deposition calculation was run at each time step using $2.5\times 10^{6}$ particles, and the Pleiades equilibrium was updated after every 100 time steps.
The results of the end plug simulations using the RealTwin model are shown in Figure 8. end plugs which failed to satisfy the DCLC constraint (33a) or in which the equilibrium failed to converge due to $\beta$ becoming too high were removed from this analysis (this is why there is an empty regions in the Figure 8 plots). A potentially surprising result of these scans is that there is little difference in peak end plug density versus $E_{\mathrm{NBI}}$, however, remembering the analysis in Section 4.1, this makes sense. Neutral beam absorption drops with higher $E_{\mathrm{NBI}}$, but the NBI fueling efficiency increases $\propto E_{\mathrm{NBI}}^{1/2}$. Another observation is that $\langle n\rangle$ improves superlinearly with increasing $P_{\mathrm{NBI}}$. This is due to a twofold non-linearity in the NBI absorption equation (33b) with NBI power. As $P_{\mathrm{NBI}}$ increases, the $\langle n\rangle$ increases and therefore the neutral beam absorption fraction increases. In addition, the plasma $\beta$ also increases as $P_{\mathrm{NBI}}$ increases. This causes the plasma to expand outwards from diagmagnetism, increasing the NBI absorption path length, further increasing NBI absorption.
Figure 9: Density $n_{p}$ and average ion energy $\langle E_{i}\rangle$ and electron energy $\langle E_{e}\rangle\equiv 1.5T_{e}$ versus the square root of the normalized poloidal flux $\sqrt{\psi_{n}}$. The profiles come from a simulation with parameters: $a_{m}=0.15~{}\mathrm{m}$, $B_{m}=25~{}\mathrm{T}$, $B_{0}=4~{}\mathrm{T}$, $\ell=4.5~{}\mathrm{m}$, $E_{\mathrm{NBI}}=240~{}\mathrm{keV}$, and $P_{\mathrm{NBI}}=15~{}\mathrm{MW}$.
As anticipated in Section 4.1, the target end plug $\langle n\rangle$ values identified in Section 3 were achieved at the lowest $P_{\mathrm{NBI}}$ for $E_{\mathrm{NBI}}$ of $240~{}\mathrm{keV}$ and $360~{}\mathrm{keV}$. Sufficient end plug conditions for fusion relevant central cell operation were achievable with $B_{0}$ of $4-6~{}\mathrm{T}$, $P_{\mathrm{NBI}}\sim 15~{}\mathrm{MW}$, $E_{\mathrm{NBI}}\sim 240-360~{}\mathrm{keV}$, and $a_{m}$ of $0.15-0.2~{}\mathrm{m}$. Under these conditions effectively full NBI absorption (>95%) was obtained with a $240~{}\mathrm{keV}$ beam, so lowering NBI energy further would only serve to reduce performance. The plasma profiles of the most optimal plug identified in this scan are shown in Figure 9. The plug had parameters: $a_{m}=0.15~{}\mathrm{m}$, $B_{0}=4~{}\mathrm{T}$, $E_{\mathrm{NBI}}=240~{}\mathrm{keV}$, and $P_{\mathrm{NBI}}=15~{}\mathrm{MW}$ and achieved $\langle n\rangle>1.5\times 10^{20}~{}\mathrm{m}^{-3}$, more than sufficient to drive a $50~{}\mathrm{m}$ long central cell to $Q>5$ conditions based on the analysis in Section 3. This plug in particular was chosen as it has a low $a_{m}$ and $E_{\mathrm{NBI}}$, simplifying engineering, while still delivering good plasma performance. Another promising, but more technologically aggressive, end plug condition that may have better DCLC stability due to its larger $a/\rho_{i}$ was obtained with $a_{m}=0.2~{}\mathrm{m}$, $B_{0}=5~{}\mathrm{T}$, $E_{\mathrm{NBI}}=360~{}\mathrm{keV}$, and $P_{\mathrm{NBI}}=20~{}\mathrm{MW}$ achieved $\langle n\rangle\sim 1.3\times 10^{20}~{}\mathrm{m}^{-3}$ also sufficient to drive a $50~{}\mathrm{m}$ central cell to $Q>5$ conditions, indicating some degree of design flexibility. The optimized tandem mirror operating points identified here are summarized in Table 3.
Parameter Optimum Alternate
End plug
$a_{m}$ 0.15 m 0.20 m

### v2

@@ 4.1 Assumptions and 0D constraints
Beginning with the modelling assumptions and fixed parameters; two large physics assumptions were made in this study to overcome the limitations of present modelling tools. First, it was assumed that the end plug, for the purposes of equilibrium and transport modelling, could be treated as a standalone simple mirror with certain correction factors. The justification of the first assumption lies in the fact that the ion populations in the plug and the central cell are to lowest order "detached". Ions confined in the central cell do not enter the plugs and ions confined in the plugs do not enter the central cell. Similarly ions that are not confined in the plug are also not confined in the central cell and vice-versa. This property of tandem mirrors ends up being useful for reasons that will be discussed momentarily. In fusion-relevant conditions, confinement is good in both the plugs and the central cell and the unconfined population of ions flowing through either region is small, acting as a (relatively) low density stream which passes outwards until it reaches the expander region (Cohen et al., 1980). However, in a classical tandem mirror the electron population is able to transit between the plug and central cell, and alpha particle heating from the central cell is sufficient to enable us to reach high electron temperatures. In fact, the predicted electron temperatures in a tandem are higher than they otherwise would be in a standalone end plug due to alpha heating of electrons in the central cell. The initial simulations in Section 4.2 of the standalone simple mirror end plug utilize a dynamic electron temperature. However, later simulations in Section 4.3 use a fixed values of $T_{e}$ and ambipolar potential $\phi_{e}$ corresponding to the values computed in the tandem POPCON.
The second important physics assumption is that MHD (Rosenbluth et al., 1957) and trapped particle mode (Berk & Lane, 1986a, b) were not considered in detail. While kinetic stability to the DCLC instability and MHD stability to the mirror and firehose instabilities are considered here, stability against flute modes and rigid $m=1$ displacements driven by MHD and trapped particle modes have not been assessed. While it is acknowledged that stability to these modes is vital to developing an operating mirror fusion device, in order to effectively assess stability and develop working stabilization actuators, self-consistent mirror equilibrium parameters, like those calculated here, must be obtained first. This work encompasses that important first step. The IPS driven RealTwin simulation framework built here was constructed for the express purpose of coupling transport, heating, and equilibrium simulations self-consistently to stability simulations. An early demonstration of this capability, in which CQL3D-m/Pleiades simulations are coupled to hybrid Vector Particle-In-Cell (hybrid-VPIC) simulations in order to assess extended MHD and kinetic stability. This is presented in this article’s companion article Tran et al. (2024). Realta Fusion plans to continue to investigate kinetic and MHD stability utilizing a full complement of simulation tools coupled to IPS driven simulations including: linear MHD (FLORA) (Cohen et al., 1986), nonlinear MHD (NIMROD) (Sovinec et al., 2004), particle-in-cell methods (VPIC, WARP-X) (Bird et al., 2021; Fedeli et al., 2022; Le et al., 2023), as well as gyrokinetic and multifluid codes (Gkeyll) (Hakim, 2008; Francisquez et al., 2023). These simulation tools can also account for extended effects: such as charge separation’s impact on trapped particle modes (Kesner & Lane, 1983), FLR effects’ impact on MHD stability (Ryutov et al., 2011), and sloshing ions as well as neutral particle fuelling’s impact on DCLC (Baldwin, 1977; Tran et al., 2024). As stability actuators and these effects can interact in ways which are nonlinear and difficult to predict analytically, high-fidelity simulation tools and self-consistent simulations are vital to developing a complete picture of stability.
This simulation model will eventually allow a systematic evaluation of axisymmetric mirror stability to trapped particle modes, MHD, and kinetic modes, using realistic equilibrium parameters to be performed under the influence of stability actuators such as: end-ring induced shear flows (Sakai et al., 1993; Beklemishev et al., 2010; Bagryansky et al., 2011), localized electron cyclotron heating induced shear flows (Cho et al., 2005, 2006), active feedback control (Lieberman & Wong, 1977; Kang et al., 1988), high-$\beta$ (Berk et al., 1987) field ripple in the central cell in combination with conformal conducting walls (Li et al., 1987), kinetic stabilizers (Post, 2001), and line-tying using cold plasma mantels (Fornaca et al., 1979; Molvik et al., 1984). In addition to MHD modes, there is interest in the impact of these actuators on trapped particle modes in the tandem mirror. Kinetic stabilizers (Post, 2001) have been shown to be potentially ineffective at stabilizing trapped particle modes (Berk & Pratt, 2011; Fowler et al., 2017). However, as trapped particle modes in tandem mirrors cause fluid-like interchange which has similar growth rates and is phenomenologically similar to MHD, it may be possible to stabilize them with the same feedback control, line tying, and shear flow actuators as MHD (Kesner, 2018). Another potential avenue for trapped particle mode stabilization is the difference in the electron and ion bounce points in the central cell which can have a stabilizing effect on trapped particle modes in classical tandem mirrors like those here (Kesner & Lane, 1983).
@@
As mentioned earlier, the detached nature of the end plug and central cell ion distributions can be used to improve performance. In a tandem, the end plug ion composition need not be the same as the central cell as confined ion populations in the end plug and central cell are not exchanged (Fowler & Logan, 1977; Cohen et al., 1980). This presents a unique opportunity to enhance end plug physics performance and simplify end plug engineering. To this end, end plugs fueled with only tritium are used in conjunction with a deuterium and tritium fueled central cell. This combination maximizes end plug performance and reduces the nuclear engineering challenges presented by the end plug. The physical argument for this operational mode is as follows: ion confinement in the end plug is dominated by two processes: ion-ion scattering and electron drag. The particle confinement times associated with these two processes are (Logan et al., 1980):
@@
At typical tandem operational parameters, confinement is assumed to be electron drag dominated, $\tau_{d}\ll\tau_{s}$ (Logan et al., 1980; Mirin et al., 1983; Hua & Fowler, 2004; Fowler et al., 2017). As the end plug ion population composition need not be the same as the central cell, performance can be enhanced by choosing ions with large $\mu/Z^{2}$. In addition, choosing an ion mix to suppress fusion neutron rates in the end plug can substantially simplify engineering by reducing radiation shielding requirements for components such as magnets and NBI. Suppressing fusion rates in the end plugs is acceptable as the total fusion power produced by the end plug is a small component of the overall fusion power in a tandem, since the end plug volume is much less than central cell volume, $V_{p}\ll V_{c}$. If taken to its logical conclusion, this argument suggests that a tritium (T) only end plug is optimal with respect to classical transport and neutron damage. Tritium has the highest $\mu/Z^{2}$ ratio of any standard fusion fuel ion, and the T-T fusion cross-section is nearly an order of magnitude lower than the D-D fusion cross section, and T-T fusion produces neutrons with energies of $\sim 3~{}\mathrm{MeV}$ (Brown et al., 2018). Thus, a T-only end plug provides a factor of $\sim 1.2$ increase in confinement performance and a many-order of magnitude reduction in the neutron rate as well as a significant reduction in neutron energy relative to a D-T end plug (T fuelled plugs will have inferior microstability to D or H plugs as the T Larmor radius is larger, making it easier to excite diamagnetic drift waves. However, it will be shown that it should be possible to avoid diamagnetic drift waves with large $a/\rho_{i}$ even with T plugs).
@@
Another assumption is the use of only negative-ion neutral beams for end plug heating in steady-state (a discussion about the limitations associated with high-harmonic RF heating coupled with positive-ion NBI, similar to what is planned in WHAM, and the reason it is not used in this work is discussed in Appendix A). The use of ECH has also been limited to achieving breakdown (assuming a scheme like that found in Yakovlev et al. (2016)). Presently, the use of any significant quantity of ECH in CQL3D-m causes the potential profile to invert. This behaviour is actually promising as it indicates that it may be possible to generate thermal barriers (Baldwin & Logan, 1979), but at the moment it causes significant problems with the bounce-averaging and ambipolar potential calculations in CQL3D-m. Thus, further investigation of ECH is left for future work. Neutral transport into the tandem mirror plasma has also been neglected. This effect is anticipated to be of some importance in small relatively low density mirrors such as WHAM, where neutrals recycled from the edge can enter the plasma and charge exchange with well-confined beam ions causing them to be lost and replacing them with a promptly lost cold ion. However, fusion-grade plasmas like those under consideration here are opaque to edge neutrals, and the effect of these neutrals on transport can be ignored. The final assumption made here is that transport in the end plug will remain classical during operation. To this end, it has been insured that the DCLC stability condition is well satisfied in the plasmas optimized here.
@@
In addition to physics assumptions, some engineering constraints were imposed on the mirror design here. First, the field of the two large bore high-field mirror coils placed at either end of the end plug (denoted HF) was fixed at $25~{}\mathrm{T}$. Feasibility studies of HTS coils at these fields have found that such designs should be viable at bores sizes appropriate for fusion applications, $\sim 1$ m warm bore diameter. A single low-field coil (denoted LF) was placed at the centre of the end plug to give control of the central field. The magnetic configuration for an example case is shown in Figure 6. The minimum length $\ell$ (where subscript $p$ on end plug quantities in this section has been dropped for simplicity), defined as the mirror throat to mirror throat distance. This limit was chosen not by the finite Larmor radius (FLR) stabilization condition for $m\geq 2$ modes, but rather by engineering constraints. The substantial size of the neutral curvature central cell in addition to the good curvature expanders is expected to provide significant stabilization against high $m$ modes (Roberts & Taylor, 1962; Rosenbluth et al., 1962; Rosenbluth, 1965; Pearlstein, 1966; Anikeev et al., 1992; Ryutov et al., 2011; Fowler et al., 2017). Active stabilization schemes are anticipated to provide for remaining MHD stabilization needs. Here, $\ell\geq 4.5~{}\mathrm{m}$ has been chosen as with values of $\ell<4.5~{}\mathrm{m}$ require large reverse polarity currents in the LF coil to cancel out the HF field. Additionally, very short end plugs will almost certainly impede NBI port access and require the use of inefficient NBI launch angles that limit beam path length, causing reduced NBI absorption. The optimal NBI configuration for maximizing beam path length, and maintaining good confinement (noting there is an effective mirror ratio reduction associated with NBI launch angles smaller than $90^{\circ}$ relative to the magnetic axis) occurs at, or very close to, a neutral beam with $\theta_{\mathrm{NBI}}=45^{\circ}$ targeted at the midplane. To maintain an effective $45^{\circ}$ ion velocity-space angle for different beam launch angles, the beam’s target location must be moved to different values of $B/B_{0}$. The reduction factor in beam path length incurred by this move is in a simple mirror is $\sqrt{(1+\cot^{2}\theta_{\mathrm{NBI}})/4}/\sin\theta_{\mathrm{NBI}}$. To simplify this study, only the near optimal case of $\theta_{\mathrm{NBI}}=45^{\circ}$ with the beam directed at the midplane was examined in detail (the NBI here used a $0.4~{}\mathrm{m}$ by $0.8~{}\mathrm{m}$ source with no beam divergence as a placeholder until more realistic NBI specifications are known). In later work, with better defined, beam, magnet, and radiation shielding parameters, this angle may be slightly adjusted to prevent beam-line interference with components.
@@
DCLC $\underline{\text{DCLC Stability}}$
$\frac{a_{0}}{\rho_{i}}\gtrsim 25$ (35b)
for $a_{m}$ and $B_{0}$. The NBI absorption condition (35a) uses the analytic NBI stopping cross sections $\sigma_{s}$ from Janev et al. (1989) and (35b) is a slightly modified version of the high-$\beta$ drift-cyclotron wave stability criterion which appears in Baldwin (1977) and Tang et al. (1972). According to these conditions, cool plasma trapped by sloshing ions or high-$\beta$ and large plasma size can be used to stabilize a plasma against DCLC. In the end plug here, stabilization of DCLC will come not from sloshing ions, but primarily from large plasma size and high-$\beta$ as the POPCONs in Section 3 predict that $T_{e}$ will be too high for effective sloshing ion trapping of cool plasma due to $\tau_{ii}\ll\tau_{d}$ (This stabilization method has been demonstrated experimentally in the LAMEX mirror (Ferron & Wong, 1984)). Using the assumptions outlined above, (35) may be rewritten (taking (35b) to be equal to 25):
@@
From the POPCON analysis performed in Section 3, it was identified that plasma radius $a_{m}\gtrsim 0.15~{}\mathrm{m}$ and density $\langle n\rangle\sim$1.0-1.5×$\times 10^{20}~{}\mathrm{m}^{-3}$. Plotting contours of $P_{\mathrm{NBI}}$ and $a_{m}$ versus $E_{\mathrm{NBI}}$ and $\langle n\rangle$ produces a plot like that shown in Figure 7. This plot indicates an operating region accessible with reasonable $P_{\mathrm{NBI}}$ values $<20~{}\mathrm{MW}$ exists between $E_{\mathrm{NBI}}$ of $200~{}\mathrm{keV}$ and $300~{}\mathrm{keV}$. However, the method used to constrain NBI absorption with (35a) is suboptimal. It has not accounted for that, despite not being absorbed as readily at high energies, the beam may still be more efficient at fuelling, and simplified scaling laws were used here rather than an actual Fokker-Planck calculation. Furthermore, the electron temperature in this analysis is implicitly taken to be $\sim 0.1E_{\mathrm{NBI}}$ in accordance with the scaling law from Egedal et al. (2022). This is lower than the values predicted by the POPCONs in Section 3. These discrepancies will be corrected in the next parts of this section with more detailed simulations using the RealTwin.
4.2 Parametric simulations of standalone simple mirror performance
Parameter Section04.2 Section04.3
$a_{m}$ 0.15 - 0.25 m 0.15 - 0.25 m